### v2

# The coisotropic embedding theorem for pre-symplectic vector spaces: an alternative proof

The coisotropic embedding theorem for pre-symplectic vector spaces: an alternative proof
L. Schiavone1,2 [Uncaptioned image]
1Dipartimento di Matematica e Applicazioni Renato Caccioppoli, Università degli Studi di Napoli Federico II
2
e-mail: luca.schiavone@unina.it
Acknowledgements
I acknowledge financial support from Next Generation EU through the project 2022XZSAFN – PRIN2022 CUP: E53D23005970006.
Introduction
The Coisotropic Embedding Theorem (CET), originally proven by Mark Gotay in his seminal work on pre-symplectic manifolds [Got82], establishes that any pre-symplectic manifold can be embedded into a symplectic manifold in such a way that the original pre-symplectic manifold becomes a coisotropic submanifold. It has significant applications both in theoretical mathematics and applied fields, where it facilitates the study of constrained mechanical systems and reduction theory from various points of view, as we are going to recall in this introduction.
In mathematical physics, Hamiltonian systems are geometrically formulated by means of symplectic manifolds (usually the cotangent bundle of a differential manifold, sometimes referred to as Phase Space) endowed with a Hamiltonian function [Arn89, AMRC78]. On the other hand, constrained Hamiltonian systems are usually described on submanifolds of the Phase Space (selected by the constraints) on which the symplectic form becomes degenerate (see [GNH77, CDI+24, CDI+22a] and references therein). Among various interesting physical systems, all gauge theories lie in this class of dynamical systems. The CET ensures that these constrained systems can still be analyzed within the framework of symplectic geometry by embedding the presymplectic manifold into a larger symplectic manifold [Got82, AMRC78, GS90].
Recent advances in mathematical physics have highlighted the utility of the CET in various contexts. For instance, in [CDI+22a], the theorem is used to establish a Noether-like one-to-one correspondence between symmetries and constants of motion within pre-symplectic Hamiltonian systems. This correspondence is fundamental in understanding the conservation laws and invariant properties of constrained mechanical systems, thereby extending the classical Noether’s theorem to a broader context. Furthermore, we employed the CET to provide a Poisson structure on the space of solutions of certain field theories. For example, in [CDI+24, CDI+22b, CDI+22c], the theorem is used to pass from a pre-symplectic framework to a symplectic one, enabling the application of symplectic geometry and Poisson structures to the field theories considered, non-Abelian gauge theories and Palatini’s formulation of General Relativity in particular. Moreover, the theorem has been instrumental in addressing the inverse problem of the calculus of variations. In recent works [Sch24b, Sch24a], we used the CET to provide a solution for the inverse problem for a class of implicit differential equations. These examples illustrate how the theorem can be leveraged to construct variational principles for systems that are not explicitly variational. Additionally, it is worth mentioning that the CET has significant applications within the Geometric Quantization program [GS81]. Geometric Quantization is a procedure for constructing a quantum theory from a classical mechanical system, and requires a symplectic structure for the classical system as starting point. By embedding a pre-symplectic manifold into a symplectic one, the theorem facilitates the Quantization of constrained systems by ensuring that the necessary symplectic structure is available.
The theorem also received several reformulations over the years. We mention the approach of V. Guilleming and S. Sternberg that proved an equivariant (with respect to the action of a Lie group) version of the theorem [GS90] and the approach of Y. G. Oh and J. S. Park [OP05] applied to the deformation of coisotropic submanifolds that makes even more transparent how the proof of the CET involves in a crucial way the use of a (arbitrary) connection providing a (arbitrary) complement of the kernel of the pre-symplectic structure.
In this work, we provide an alternative, even if equivalent, proof of the theorem. We limit ourselves to the context of vector spaces.
The content of this work is organized as follows.
In Section 1, we recall Oh and Park’s proof of the CET for pre-symplectic vector spaces.
In Section 2, we introduce our alternative proof of the theorem for pre-symplectic vector spaces.
1 Classical version of the Coisotropic Embedding Theorem
In this section, we recall the classical version of the CET for pre-symplectic vector spaces.
Let us first start with some useful definitions and results.
Definition 1.1 (Symplectic vector space).
A symplectic vector space is a vector space $V$ equipped with a skew-symmetric, bilinear, non-degenerate form $\omega$, say $(V,\,\omega)$.
Here, we recall Darboux theorem for vector spaces which is a corollary of a more general result available for smooth differential manifolds (see [AMR88]).
Theorem 1.2 (Darboux theorem for symplectic vector spaces).
Given a symplectic vector space $(V,\,\omega)$, there exists a basis:
=$\mathcal{B}\,=\,\left\{\,\frac{\partial}{\partial x_{1}},\,...,\,\frac{\partial}{\partial x_{m}},\,\frac{\partial}{\partial y^{1}},\,...,\,\frac{\partial}{\partial y^{m}}\,\right\}\,,$
[…]
$\left\{\,x^{1},\,...,\,x^{m},\,y_{1},\,...,\,y_{m}\,\right\}\,,$
such that, given its dual basis:
=$\mathcal{B}^{\star}\,=\,\left\{\,{\rm d}x_{1},\,...,\,{\rm d}x_{m},\,{\rm d}y^{1},\,...,\,{\rm d}y^{m}\,\right\}\,,$
the symplectic form $\omega$ reads:
$\omega\,=\,{\rm d}x_{j}\wedge{\rm d}y^{j}\,,$ (1)
where $\wedge$ is the skew-symmetric tensor product of elements in $\mathcal{B}^{\star}$.
Definition 1.3 (Symplectic orthogonal of a vector subspace).
Consider a symplectic vector space $(V,\,\omega)$. Consider a vector subspace $W$ of $V$. The symplectic orthogonal of $W$, say $W^{\perp}$ is defined as the set of vectors $v\in V$ such that:
[…]
Definition 1.4 (Coisotropic subspace of a symplectic vector space).
Given a symplectic vector space $(V,\,\omega)$, a vector subspace $W$ of $V$ is said to be coisotropic if $W^{\perp}\subseteq W$.
Definition 1.5 (Pre-symplectic vector space).
A pre-symplectic vector space is a vector space $V$ equipped with a skew-symmetric, bilinear form $\omega$, say $(V,\,\omega)$.
Thus, $\omega$ here may be degenerate. We will always assume here $\mathrm{rank}\,\omega\,=\,r<n$ to be constant.
Here, we recall the analogous of Theorem 1.2 for pre-symplectic vector spaces. Again, this is a corollary of a more general result available for smooth differential manifolds [AMR88].
Theorem 1.6 (Darboux theorem for pre-symplectic vector spaces).
Given a pre-symplectic vector space $(V,\,\omega)$, there exists a basis:
=$\mathcal{B}\,=\,\left\{\,\frac{\partial}{\partial x_{1}},\,...,\,\frac{\partial}{\partial x_{m}},\frac{\partial}{\partial y^{1}}\,...,\,\frac{\partial}{\partial y^{m}},\,\frac{\partial}{\partial z^{1}},\,...,\,\frac{\partial}{\partial z^{r}}\,\right\}\,,$
where $n=2m+r$ is the dimension of $V$, $r$ being the dimension of $\mathrm{ker}\,\omega$, with the associated system of coordinates on $V$:
$\left\{\,x_{1},\,...,\,x_{m},\,y^{1},\,...,\,y^{m},\,z^{1},\,...,\,z^{r}\,\right\}\,,$
such that, given its dual basis:
=$\mathcal{B}^{\star}\,=\,\left\{\,{\rm d}x_{1},\,...,\,{\rm d}x_{m},\,{\rm d}y^{1},\,...,\,{\rm d}y^{m},\,{\rm d}z^{1},\,...,\,{\rm d}z^{r}\,\right\}\,,$
the pre-symplectic form $\omega$ reads:
$\omega\,=\,{\rm d}x_{j}\wedge{\rm d}y^{j}\,,$ (3)
and, thus, $\mathrm{ker}\,\omega$ is spanned, at each point of $V$, by:
$\mathrm{ker}\,\omega\,=\,\langle\left\{\,\frac{\partial}{\partial z^{1}},\,...,\,\frac{\partial}{\partial z^{r}}\,\right\}\rangle\,.$ (4)
Let us now recall the classical way of proving the coisotropic embedding theorem.
Theorem 1.7 (Coisotropic embedding theorem for vector spaces).
Let $(V,\omega)$ be a pre-symplectic vector space. There exists a symplectic vector space $(\tilde{V},\tilde{\omega})$ such that $(V,\,\omega)$ is a coisotropic subspace of $(\tilde{V},\,\tilde{\omega})$. The vector space $(\tilde{V},\,\tilde{\omega})$ is called the symplectic thickening of $(V,\,\omega)$.
The proof of the theorem follows three main steps:
  • Construction of the symplectic thickening. Consider a pre-symplectic vector space $(V,\omega)$. Denote by $K$ the kernel of $\omega$, which is by assumption the same for each point of $V$. The vector space $V\times K$ is a vector subspace of $\mathbf{T}V\simeq V\times V$ (where $\simeq$ denote that the two spaces are isomorphic). Consider the cotangent bundle $\mathbf{T}^{\star}V\,\simeq V\times V^{\star}$, where $V^{\star}$ denotes the dual vector space of $V$. Consider the dual of $K$, say $K^{\star}$. Then, the symplectic thickening of $V$ is the vector space $\tilde{V}\,=\,V\times K^{\star}$ which is a vector subspace of $\mathbf{T}^{\star}V$. In the next step we prove why such a symplectic thickening is a symplectic vector space.
  • Construction of the symplectic structure. Consider a complement $W$ of $K$ into $V$, namely a vector subspace of $V$ such that $V\,=\,W\oplus K$. It is not unique. Indeed, since $K$ is spanned by:
    $\left\{\,\frac{\partial}{\partial z^{1}},\,...,\,\frac{\partial}{\partial z^{r}}\,\right\}\,,$
    the complement $W$ may be any of the following vector subspaces of $V$:
    $W\,=\,\langle\,\left(\,\frac{\partial}{\partial x_{j}}+{P_{x}}^{ja}\frac{\partial}{\partial z^{a}}\,\right)_{j=1,...,m;a=1,...,r},\,\left(\,\frac{\partial}{\partial y^{j}}+{P_{y}}_{j}^{a}\frac{\partial}{\partial z^{a}}\,\right)_{j=1,...,m;a=1,...,r}\,\rangle\,,$ (5)
    where ${P_{x}}^{ja}$ and ${P_{y}}_{j}^{a}$ are the entries of rectangular matrices depending on $(x,\,y,\,z)$. The vectors appearing in the latter equation, which are parametrized by the arbitrary values of ${P_{x}}^{ja}$ and ${P_{y}}_{j}^{a}$, span the kernel of the following system of linear forms on $V$:
    $P^{a}\,=\,{\rm d}z^{a}-{P_{x}}^{ja}{\rm d}x_{j}-{P_{y}}^{a}_{j}{\rm d}y^{j}\,.$ (6)
    Such a system of linear forms fulfill the conditions:
    $P^{b}\left(\,\frac{\partial}{\partial z^{a}}\,\right)\,=\,\delta_{a}^{b}\,,$ (7)
    which ensures that the (1, 1)11(1,\,1)( 1 , 1 )-tensor:
    $P\,:=\,P^{a}\otimes\frac{\partial}{\partial z^{a}}\,,$ (8)
    is idempotent, namely it is a "projector" onto $K$. Therefore, the (1, 1)11(1,\,1)( 1 , 1 )-tensor just constructed is such that its image is $K$ and its kernel is the complementary space $W$. Such a tensor is what is usually referred to as a connection on $V$ and it is a useful way of parametrizing the arbitrariness in choosing the complement $W$.
    The decomposition $V\,=\,W\oplus K$ provided by $P$ allows us to extend $\omega$ to a symplectic form on $\tilde{V}$. Indeed, associated to $P$, there is the following linear form on $\tilde{V}$:
    $\theta^{P}\,=\,{p_{z}}_{a}P^{a}\,,$ (9)
    where $\left\{\,{p_{z}}_{a}\,\right\}_{a=1,...,r}$ is a system of coordinates on $K^{\star}$. The linear form $\theta^{P}$ evidently depends on the particular connection $P$ chosen as stressed by the notation. It allows one to define a symplectic form on $\tilde{V}$ as:
    $\tilde{\omega}\,=\,\tau^{\star}\omega+{\rm d}\theta^{P}\,,$ (10)
    where $\tau$ is the projection from $V\times K^{\star}$ to $V$. It reads:
    $\tilde{\omega}\,=\,\tau^{\star}\omega+{\rm d}{p_{z}}_{a}\wedge P^{a}+{p_{z}}_{a}{\rm d}P^{a}\,.$ (11)
    This form is symplectic. Indeed, it is skew-symmetric and bilinear by definition and it has no kernel as it can be proven via a straightforward computation.
  • Check of the coisotropic condition. The original pre-symplectic vector space $V$ can be naturally embedded as the zero-section of $\tau$ into $\tilde{V}$, namely, it can be recovered as the subspace of $\tilde{V}$ selected by the conditions
[…]
    A direct computation shows that the symplectic orthogonal of $V$ with respect to the symplectic structure $\tilde{\omega}$ is void, namely $V$ is trivially a coisotropic subspace of $\tilde{V}$.
2 A new proof of the Coisotropic Embedding Theorem
Here, we propose our alternative proof of the theorem.
Our proof will follow the same three steps as the classical one:
  • Construction of the symplectic thickening. Given a pre-symplectic vector space $(V,\,\omega)$, the cotangent bundle $(\mathbf{T}^{\star}V,\,\omega_{\mathbf{T}^{\star}V})\,\simeq\,(V\times V^{\star},\omega_{\mathbf{T}^{\star}V})$ is a symplectic vector space, equipped with the canonical symplectic structure $\omega_{\mathbf{T}^{\star}V}$. Denote by $\rho$ the canonical projection $\rho\;\colon\;\;\mathbf{T}^{\star}V\to V$. Denote by:
    $\{\,x_{j},\,y^{j},\,z^{a}\,\}_{j=1,...,m;a=1,...,r}$
    ($r$ being again the dimension of $\mathrm{ker}\,\omega$ and $n=2m+r$ being the dimension of $V$) a system of Darboux coordinates on $V$ and by:
    $\{\,x_{j},\,y^{j},\,z^{a},\,{p_{x}}^{j},\,{p_{y}}_{j},\,{p_{z}}_{a}\,\}_{j=1,...,m;a=1,...,r}$
    an adapted system of coordinates on $\mathbf{T}^{\star}V$. In this system of coordinates
    $\omega\,=\,{\rm d}x_{j}\wedge{\rm d}y^{j}\,,$ (12)
    and
    $\omega_{\mathbf{T}^{\star}M}\,=\,{\rm d}{p_{x}}^{j}\wedge{\rm d}x_{j}+{\rm d}{p_{y}}_{j}\wedge{\rm d}y^{j}+{\rm d}{p_{z}}_{a}\wedge{\rm d}z^{a}\,.$ (13)
    The form $\omega^{\prime}\,=\,\omega_{\mathbf{T}^{\star}V}+\tau^{\star}\omega$ is symplectic as well. Indeed, to prove that it is non-degenerate it is sufficient to prove that
    $\omega^{\prime n}\,=\,\omega^{\prime}\underbrace{\,\wedge\,...\,\wedge\,}_{n\,\text{times}}\omega^{\prime}$
    is different from zero for $n$ equal to half the dimension of $\mathbf{T}^{\star}V$. A direct computation shows that $\omega^{\prime n}\,=\,\omega_{\mathbf{T}^{\star}V}^{n}$, where $\omega_{\mathbf{T}^{\star}V}^{n}$ is different from zero because $\omega_{\mathbf{T}^{\star}V}$ is symplectic.
    In the system of coordinates chosen, $\omega^{\prime}$ reads
    $\omega^{\prime}\,=\,{\rm d}{p_{x}}^{j}\wedge{\rm d}x_{j}+{\rm d}{p_{y}}_{j}\wedge{\rm d}y^{j}+{\rm d}{p_{z}}_{a}\wedge{\rm d}z^{a}+{\rm d}x_{j}\wedge{\rm d}y^{j}\,.$ (14)
    Recall that $\left\{\,\frac{\partial}{\partial z^{a}}\,\right\}_{a=1,...,r}$ is a system of generators of $\mathrm{ker}\,\omega\,=\,K$. Their cotangent lifts are those vector fields $\tilde{K}_{a}$ tangent to $\mathbf{T}^{\star}V$ defined by the condition
[…]
    where $\mathcal{L}_{\tilde{K}_{a}}$ denotes the Lie derivative along the vector field $\tilde{K}_{a}$. They are computed to be
    $\tilde{K}_{a}\,=\,\frac{\partial}{\partial z^{a}}\,.$ (16)
    Denote by $\left\{\,H_{a}\,\right\}_{a=1,...,r}$ the Hamiltonian functions associated to $\tilde{K}_{a}$ via $\omega^{\prime}$, namely those functions staisfying
    $i_{\tilde{K}_{a}}\omega^{\prime}\,=\,{\rm d}H_{a}\,,$ (17)
    where $i_{\tilde{K}_{a}}$ denotes the contraction of the differential form along the vector field $\tilde{K}_{a}$. They read
    $H_{a}\,=\,{p_{z}}_{a}\,.$ (18)
    These Hamiltonian functions, together with the vectors in the original pre-symplectic vector space $V$, determine a subspace of $\mathbf{T}^{\star}V$
    $\tilde{V}\,=\,\left\{\,\frac{\partial}{\partial x_{j}},\,\frac{\partial}{\partial y^{j}},\,\frac{\partial}{\partial z^{a}},\,\frac{\partial}{\partial{p_{z}}_{a}}\,\right\}_{j=1,...,m,\,a=1,...,r}\,,$ (19)
    namely the same constructed in Section 1.
    There are many, non-equivalent ways of embedding $\tilde{V}$ into $\mathbf{T}^{\star}V$. In general, an embedding of $\tilde{V}$ into $\mathbf{T}^{\star}V$ is any map of the type:
    $\begin{split}\tilde{V}\to\mathbf{T}^{\star}V\;\;:\;\;&(\,x_{j},\,y^{j},\,z^{a},\,{p_{z}}_{a}\,)\\ \mapsto&\left(\,x_{j},\,y^{j},\,z^{a},\,{p_{x}}^{j}\,=\,f^{j}(x,\,y,\,z,\,p_{z}),\,{p_{y}}_{j}\,=\,g_{j}(x,\,y,\,z,\,p_{z}),\,{p_{z}}_{a}\,\right)\,,\end{split}$ (20)
    with $f^{j}$ and $g_{j}$ smooth functions of their arguments. However, here we are going to consider only those which preserve the cotangent structure of $\mathbf{T}^{\star}V$. This class of embeddings is of the type
    $\begin{split}\mathfrak{i}\;\colon\;\;\tilde{V}\to\mathbf{T}^{\star}V\;\;:\;\;&(\,x_{j},\,y^{j},\,z^{a},\,{p_{z}}_{a}\,)\\ \mapsto&\left(\,x_{j},\,y^{j},\,z^{a},\,{p_{x}}^{j}\,=\,-{P_{x}}^{ja}{p_{z}}_{a},\,{p_{y}}_{j}\,=\,-{P_{y}}^{a}_{j}{p_{z}}_{a},\,{p_{z}}_{a}\,\right)\,,\end{split}$ (21)
    where ${P_{x}}^{ja}$ and ${P_{y}}^{a}_{j}$ are the entries of two rectangular matrices of maximal rank depending on $(x,\,y,\,z)$ and where the minus sign is an arbitrary choice. The fact that $\mathfrak{i}$ is linear in the ${p_{z}}_{a}$ ensures that $\mathfrak{i}$ preserves the zero-section of $\rho$, namely $\rho\bigr{|}_{Im\,\mathfrak{i}}$ is still a submersion onto $V$. On the other hand, the maximal rank hypothesis ensures that the fibres of $\rho\bigr{|}_{Im\,\mathfrak{i}}$ are all diffeomorphic, and, thus, the image of $\mathfrak{i}$ defines a vector subbundle of $\mathbf{T}^{\star}V$.
  • Construction of the symplectic structure. The symplectic structure on $\tilde{V}$ here is constructed by simply restricting $\omega^{\prime}$ to $\tilde{V}$ by taking its pull-back along $\mathfrak{i}$. Such a pull-back reads:
    $\begin{split}\mathfrak{i}^{\star}\omega^{\prime}\,=:\,\tilde{\omega}\,=\,&dx_{j}\wedge dy^{j}+\\ +&d{p_{z}}_{a}\wedge\left(\,{\rm d}z^{a}-{P_{x}}^{ja}{\rm d}x_{j}-{P_{y}}_{j}^{a}{\rm d}y^{j}\,\right)+\\ -&{p_{z}}_{a}\left(\,{\rm d}{P_{x}}^{ja}\wedge{\rm d}x_{j}+{\rm d}{P_{y}}_{j}^{a}\wedge{\rm d}y^{j}\,\right)\,,\end{split}$ (22)
    that coincides with (11). One can check that $\tilde{\omega}$ is non-degenerate only within the intersection of the image of $\mathfrak{i}$ and of the zero-section of $\mathbf{T}^{\star}V$.
  • Check of the coisotropic condition. To prove that $V$ is a coisotropic vector subspace of $(\tilde{V},\,\tilde{\omega})$, it is sufficient to compute $V^{\perp}$ and to note that it is made only by the zero vector. Consequently $V^{\perp}\subseteq V$ and, thus, $V$ is coisotropic in $(\tilde{V},\,\tilde{\omega})$.
Conclusions
We provided an alternative proof of the CET for pre-symplectic vector spaces that passes through the cotangent bundle of the pre-symplectic vector space and individuates the symplectic thickening as the subspace of such a cotangent bundle selected by the Hamiltonian functions associated with the vectors in the kernel of the original pre-symplectic form. The classical ambiguity in choosing a complement for the kernel of the pre-symplectic structure is recovered as the ambiguity in embedding such a subspace into the cotangent bundle over the original pre-symplectic vector space.